# On a conjecture concerning the Fisher–Widom line and the line of vanishing excess isothermal compressibility in simple fluids

### v1

@@ 1 Introduction
Another example related to the thermodynamic properties is the compressibility factor defined as $Z(\rho,T)={p}/{\rho k_{B}T}$, where $p$ is the pressure, $\rho$ is the number density, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature. As is well known, $Z=1$ for an ideal gas. When the attractive part of the potential dominates (low enough temperatures and/or densities), then $Z$ tends to be smaller than $1$, while if the repulsive part dominates (high enough temperatures and/or densities, $Z$ tends to be greater than $1$. In the phase diagram of a simple fluid, the line in the temperature vs density plane separating the region where $Z<1$ from the one in which $Z>1$ is called the Zeno line [22]. It is assumed to be an almost straight line that starts at the Boyle temperature and ends by crossing the density axis at the so-called Boyle density $\rho_{B}$, which is the value of the density obtained by extrapolating the coexistence curve into the low-temperature region beyond the triple point. The value of another thermodynamic quantity, the isothermal susceptibility $\chi_{T}(\rho,T)=k_{B}T\left({\partial\rho}/{\partial p}\right)_{T}$, which is equal to $1$ for an ideal gas, also serves to indicate whether it is the attractive part of the potential the one that dominates (when $\chi_{T}>1$) or whether the repulsive part is the dominant one (when $\chi_{T}<1$). The line in the phase diagram with $\chi_{T}=1$ (which also starts at the Boyle temperature in the temperature vs density plane) separates the regions where either part of the potential dominates from the perspective of the isothermal susceptibility. The line $\chi_{T}=1$ has been referred to in the literature as the ‘line of vanishing excess isothermal susceptibility’ [17]. However, in analogy with the reasoning [22] that led to coin the term ‘Zeno’ line ($Z=1$), from here onwards, and for reasons to be explained below, we will abbreviate the nomenclature and refer to the line $\chi_{T}=1$ as the ‘Seno’ line.
The above discussion has focused on qualitative arguments related to (in principle) measurable thermodynamic quantities. We now turn specifically to structural properties. The statistical mechanics expression for the compressibility factor in $d$ dimensions, as obtained from the virial route, gives $Z$ in terms of the intermolecular potential $\phi(r)$ and the radial distribution function $g(r)$ as [23, 24]
@@
The role played by the attractive and repulsive parts of the potential on the structural properties of simple fluids is best exemplified by the study (first carried out by Fisher and Widom [1] for one-dimensional lattice-continuum models) of the asymptotic decay of the total correlation function. In fact, the effect of a dominant repulsive part manifests itself in a damped oscillatory decay, while the decay is monotonic if the dominant part is the attractive one. The so-called Fisher–Widom (FW) line in the temperature vs density plane of the phase diagram is the line that separates these two regions, namely the region in which the asymptotic decay of $h(r)$ is monotonic and the region in which it is damped oscillatory. This line has received a lot of attention and, recently, Stopper et al. [17] have conjectured that the Seno line should approximate well the FW line in simple fluids. It is the main aim of this paper to examine such a conjecture by considering a particular model potential, the Jagla potential [25] (hard core plus a linear repulsive ramp and a linear attractive ramp) given by
@@
This paper was prepared as an invited contribution to a special issue of Molecular Physics in honor of Luis F. Rull and José Luis Fernández Abascal. Apart from the fact that Luis addressed the problem of the location of the FW line for systems interacting through short-ranged potentials [5] and so our contribution is clearly aligned with the purpose of the special issue, we want to stress the personal connection of Luis with two of us (A.S. and A.R.R.). In this regard, we should mention that the first scientific paper that A.S. published [26] involved a collaboration with him. On the other hand, Luis was also the head of the group in which A.R.R. carried out his Ph. D. thesis and together with Luis he published three papers [27, 28, 29].
The paper is organised as follows. In Section 2, we present the calculations pertaining to the one-dimensional Jagla fluid (in which case exact results may be derived) for the Zeno line, the Seno line and the FW line. This is followed in Section  3 by parallel calculations for the three-dimensional system, where we have used the theoretical rational-function approximation [30, 31, 24] and Monte Carlo computer simulations. The paper is closed in Section 4 with a discussion of the results and some concluding remarks. Some mathematical details have been relegated to an Appendix.
@@
For the FW line, one needs the nonzero poles of $G(s)$, i.e., the roots of the equation $\Omega(s+\beta p,\beta)=\Omega(\beta p,\beta)$, with the least negative real part, since these will determine the asymptotic behaviour of the total correlation function $h(r)$. Near the FW line, the dominant poles are either a pair of complex conjugates ($s=-\zeta\pm i\omega$) or a real value ($s=-\kappa$), so that
@@
where $\delta$ is the argument of the residue $A_{\zeta}$ , i.e., $A_{\zeta}=|A_{\zeta}|e^{\pm i\delta}$ and $\kappa^{-1}$ is the correlation length. Once the poles have been computed, the FW line may readily be obtained as the locus of points where $\zeta=\kappa$, that is
@@
Up to this point, we now have all the necessary ingredients to compute the Zeno, Seno and FW lines for the one-dimensional Jagla fluid. But before doing that, and for the sake of completeness, we will take advantage of the simple form of the intermolecular potential $\phi(r)$ of this fluid, as given by Equation (3), to obtain explicitly its second virial coefficient. This will provide us with the means to compute also the Boyle temperature. The explicit analytic result for the second virial coefficient reads
@@
Figure 1: Zeno, Seno and FW curves in the $T^{*}$ vs $\rho^{*}$ plane for a one-dimensional Jagla fluid with the parameters given in Equation (6). The open circle at $\rho^{*}=0$ represents the Boyle temperature $T_{B}^{*}\simeq 0.4758$. Below the Zeno line, one has $Z<1$, while $Z>1$ above it. Similarly, $\chi_{T}>1$ below the Seno line and $\chi_{T}<1$ above it. Furthermore, below the FW line, the decay of $h(r)$ is monotonic, while it is oscillatory above it.
In Figure 1 we show the resulting Zeno, Seno and FW lines in the temperature vs density plane. Note that, while the Zeno and Seno lines start at the Boyle temperature, the FW line diverges for $\rho\to 0$. Two more things are also worth pointing out at this stage. On the one hand, the Zeno line is not a straight line and all three curves end at the Boyle density $\rho_{B}=\lambda_{1}^{-1}$ (see the Appendix for a proof). On the other hand, it is clear that for this system the conjecture of Stopper et al. [17] concerning the Seno and FW lines is not sustained. Whether it will hold for the three-dimensional Jagla fluid will be discussed in Section 3.
@@ 3.2 Rational-function approximation
In a previous paper [16] some of us presented a semi-analytical approach based on the rational-function approximation (RFA) [30, 31, 24] to obtain $g(r)$, including its asymptotic behaviour for large $r$. The application of the RFA to the Jagla fluid was made by assuming that a discretised version of the potential given in Equation (3) consisting in a hard core plus of a sequence of $n$ steps of heights $\varepsilon_{j}$ and widths $\sigma_{j}-\sigma_{j-1}$ (with the conventions $\sigma_{0}=\sigma$ and $\sigma_{n}=\lambda_{2}$), leads to essentially the same cavity function as the original Jagla potential. By considering the second virial coefficient and some representative cases, it was found that the choice $n=10$ proved to be a reasonable one, leading to good agreement with MC simulation results. Such an agreement worsened as the density increased and/or the temperature decreased, especially near contact. But, even in those cases, the oscillations of $g(r)$ for larger distances were well accounted for, at least at a qualitative level.
@@ 3.3 Monte Carlo simulations
For the FW line, the main problem is how to know from the MC data of $g(r)$ at a given state $(\rho^{*},T^{*})$ sufficiently close to the line whether that state is above the line (region of oscillatory decay) or below it (region of monotonic decay). If $g(r)$ were known with a good signal-to-noise ratio in the asymptotic large-$r$ domain, it would be in principle possible to assess whether the decay is oscillatory or monotonic since one of the two competing behaviours in Equation (17) would dominate. However, the closer the state is to the line, the closer the values of $\kappa$ and $\zeta$ become. Consequently, larger distances are required to observe the prevalence of one of the two competing behaviours. In addition, it is worth noting that the amplitude $A_{\kappa}$ of the monotonic behaviour is typically smaller than the amplitude $2|A_{\zeta}|$ of the oscillatory behaviour. As a result, the oscillatory behaviour can overshadow the monotonic behaviour for intermediate distances, even if $\kappa<\zeta$. We have also noted that the fitting in Equation (19), although suitable for measuring $\chi_{T}$, lacks robustness in determining whether $\kappa<\zeta$ or $\kappa>\zeta$.
@@
Figure 2 shows $r|h(r)|$ (in logarithmic scale), as obtained from the RFA, for a density $\rho^{*}=0.30$ and four temperatures: $T^{*}=0.60$, $0.65$, $0.70$ and $0.75$. For this density, the RFA temperature corresponding to the FW line is known to be $T^{*}=0.7315$. The first thing to note in this case is that the approximation with the two leading poles is able to capture the whole total correlation function for distances beyond $r\simeq 3\sigma$ for all temperatures. Next, we note that the signature that one is sufficiently below the temperature corresponding to the FW line is that ‘anomalous’ neighbouring nodes appear [cf. Figures 2(a–c)]. These nodes, which eventually disappear for large enough distances [although oscillations may still be seen, cf. Figure 2(a)], exhibit an anomalous behavior: their separation is smaller than that of neighbouring nodes and their maxima always fall below that of the neighbouring peaks [cf. Figure 2(a)]. As the temperature increases, remaining below the FW line, the anomalous nodes become progressively less apparent within the range $r<10\sigma$ [cf. Figures 2(b, c)]. Finally, when one is close to or above the temperature corresponding to the FW line the nodes become regular [cf. Figure 2(d)]. According to our criterion, one would conclude that $T^{*}=0.65$, or even $T^{*}=0.70$, are lower-bound estimates for the temperature of the FW line when $\rho^{*}=0.30$, which agrees with the true FW temperature $T^{*}=0.7315$ predicted by the RFA for $\rho^{*}=0.30$.
@@
Figure 4: Zeno, Seno and FW curves in the $T^{*}$ vs $\rho^{*}$ plane for a three-dimensional Jagla fluid with the parameters given in Equation (6). The lines are RFA predictions and the symbols represent estimates obtained from our MC simulations. The open circle at $\rho^{*}=0$ represents the Boyle temperature $T_{B}^{*}\simeq 1.3879$. Below the Zeno line, one has $Z<1$, while $Z>1$ above it. Similarly, $\chi_{T}>1$ below the Seno line and $\chi_{T}<1$ above it. Furthermore, below the FW line, the decay of $h(r)$ is monotonic, while it is oscillatory above it.
@@ 3.4 Results
In order to set the proper perspective for the assessment of our findings, in Figure 4 we show the resulting Zeno, Seno and FW lines for the three-dimensional Jagla fluid, as obtained both from the RFA approach and from simulation.
@@
Figure 5: Plot of $g(r)$ for a three-dimensional Jagla fluid with the parameters given in Equation (6) and $\rho^{*}=0.30$, $T^{*}=0.75$. The solid line corresponds to the RFA, while the dashed line represents MC simulation data. Note that the hard-core diameter $\sigma=1$ has been taken as the unit of length.
To illustrate how the discrepancies between the RFA and MC simulations for the transition lines are consistent with a reasonable global agreement in the radial distribution function, we compare the RFA and MC values of $g(r)$ at a density $\rho^{*}=0.30$ and a temperature $T^{*}=0.75$ in Figure 5. These conditions correspond to the scenarios depicted in Figures 2(d) and 3(a), respectively. A remarkable overall agreement is observed, although the RFA tends to slightly underestimate $g(r)$ within the interval $\sigma\leq r\leq\lambda_{1}$ and near the second maximum. Considering Equations (2) and (15), this suggests that the RFA tends to underestimate the values of $Z$ and $\chi_{T}$. Consequently, this leads to an upward shift of the Zeno line and a downward shift of the Seno line with respect to the MC values. It is also evident from Figure 5 that $g(r)\simeq 1$ for $r>4\sigma$ if $\rho^{*}=0.30$ and $T^{*}=0.75$. This makes it rather challenging to determine whether the asymptotic decay is monotonic or oscillatory. In the case of the RFA, we know from the pole analysis of the Laplace transform $\mathcal{G}(s)$ that the decay is oscillatory, while our criterion suggests that the decay of the MC data is monotonic [cf. Figure 3(a)].

### v2

@@ 1 Introduction
Another example related to the thermodynamic properties is the compressibility factor defined as $Z(\rho,T)={p}/{\rho k_{B}T}$, where $p$ is the pressure, $\rho$ is the number density, $k_{B}$ is the Boltzmann constant and $T$ is the absolute temperature. As is well known, $Z=1$ for an ideal gas. When the attractive part of the potential dominates (low enough temperatures and/or densities), then $Z$ tends to be smaller than $1$, while if the repulsive part dominates (high enough temperatures and/or densities), $Z$ tends to be greater than $1$. In the phase diagram of a simple fluid, the line in the temperature vs density plane separating the region where $Z<1$ from the one in which $Z>1$ is called the Zeno line [22]. It is generally assumed to be an almost straight line that starts at the Boyle temperature and ends by crossing the density axis at the so-called Boyle density $\rho_{B}$, which is the value of the density obtained by extrapolating the coexistence curve into the low-temperature region beyond the triple point. However, very recently Paterson et al. [23] have found that, for both attractive square-well fluids with varying well-widths and Mie $n$-6 fluids with different repulsive exponents $n$, irrespective of the values of the well-width or of the repulsive exponent, the corresponding Zeno lines are curved. We will come back to this point later on.
The value of another thermodynamic quantity, the isothermal susceptibility (or reduced isothermal compressibility) $\chi_{T}(\rho,T)=k_{B}T\left({\partial\rho}/{\partial p}\right)_{T}$, which is equal to $1$ for an ideal gas, also serves to indicate whether it is the attractive part of the potential the one that dominates (when $\chi_{T}>1$) or whether the repulsive part is the dominant one (when $\chi_{T}<1$). The line in the phase diagram with $\chi_{T}=1$ (which also starts at the Boyle temperature in the temperature vs density plane) separates the regions where either part of the potential dominates from the perspective of the isothermal compressibility. The line $\chi_{T}=1$ has been referred to in the literature as the ‘line of vanishing excess isothermal compressibility’ [17]. However, in analogy with the reasoning [22] that led to coin the term ‘Zeno’ line ($Z=1$), from here onwards, and for reasons to be explained below, we will abbreviate the nomenclature and refer to the line $\chi_{T}=1$ as the ‘Seno’ line.
The above discussion has focussed on qualitative arguments related to (in principle) measurable thermodynamic quantities. We now turn specifically to structural properties. The statistical mechanics expression for the compressibility factor in $d$ dimensions, as obtained from the virial route, gives $Z$ in terms of the intermolecular potential $\phi(r)$ and the radial distribution function $g(r)$ as [24, 25]
@@
The role played by the attractive and repulsive parts of the potential on the structural properties of simple fluids is best exemplified by the study (first carried out by Fisher and Widom [1] for one-dimensional lattice-continuum models) of the asymptotic decay of the total correlation function. In fact, the effect of a dominant repulsive part manifests itself in a damped oscillatory decay, while the decay is monotonic if the dominant part is the attractive one. The so-called Fisher–Widom (FW) line in the temperature vs density plane of the phase diagram is the line that separates these two regions, namely the region in which the asymptotic decay of $h(r)$ is monotonic and the region in which it is damped oscillatory.
Although not directly linked to the dominance of the attractive or repulsive part of the potential, but rather to liquid-like behaviour in the supercritical region, there is another interesting line in the temperature-density plane, the so-called Widom line [11]. This line is the locus of points of maximal response (for instance, maximal correlation length) for each temperature. As temperature decreases, the Widom line ends at the critical point, thus representing an extension of the coexistence line into the one-phase region.
The FW line has received a lot of attention and, recently, Stopper et al. [17] have conjectured that the Seno line should approximate well the FW line in simple fluids. They tested their hypothesis in a few models (square-well, hard-core Yukawa, sticky hard spheres and Asakura–Oosawa) and also located various lines relative to the gas-liquid phase coexistence, as well as the Widom line. It is the main aim of this paper to examine Stopper et al.’s conjecture by considering a particular model potential, the Jagla potential [26] (hard core plus a linear repulsive ramp and a linear attractive ramp) given by
@@
This choice for the values of the parameters is motivated by the fact that in the one-dimensional case the exact results require a nearest-neighbour interaction. On the other hand, for such values the three-dimensional Jagla fluid does not show a liquid-liquid phase separation [27].
This paper was prepared as an invited contribution to a special issue of Molecular Physics in honor of Luis F. Rull and José Luis Fernández Abascal. Apart from the fact that Luis addressed the problem of the location of the FW line for systems interacting through short-ranged potentials [5] and so our contribution is clearly aligned with the purpose of the special issue, we want to stress the personal connection of Luis with two of us (A.S. and A.R.R.). In this regard, we should mention that the first scientific paper that A.S. published [28] involved a collaboration with him. On the other hand, Luis was also the head of the group in which A.R.R. carried out his Ph. D. thesis and together with Luis he published three papers [29, 30, 31], which gave him the opportunity to start his career as a researcher in the statistical physics of liquids.
The paper is organised as follows. In Section 2, we present the calculations pertaining to the one-dimensional Jagla fluid (in which case exact results may be derived) for the Zeno, Seno, FW and Widom lines. This is followed in Section  3 by parallel calculations for the three-dimensional system, where we have used the theoretical rational-function approximation (RFA) [32, 33, 25] and Monte Carlo (MC) computer simulations. The paper is closed in Section 4 with a discussion of the results and some concluding remarks. Some mathematical details have been relegated to an Appendix.
@@
For the FW line, one needs the nonzero poles of $G(s)$, i.e. the roots of the equation $\Omega(s+\beta p,\beta)=\Omega(\beta p,\beta)$, with the least negative real part, since these will determine the asymptotic behaviour of the total correlation function $h(r)$. Near the FW line, the dominant poles are either a pair of complex conjugates ($s=-\zeta\pm i\omega$) or a real value ($s=-\kappa$), so that
@@
where $\delta$ is the argument of the residue $A_{\zeta}$ , i.e. $A_{\zeta}=|A_{\zeta}|e^{\pm i\delta}$ and $\kappa^{-1}$ is the correlation length. Once the poles have been computed, the FW line may readily be obtained as the locus of points where $\zeta=\kappa$, that is
@@
As for the Widom line, it is obtained from Equation (11c), together with the condition $(\partial\kappa/\partial\beta p)_{\beta}=0$. Deriving both sides of Equation (11c) with respect to $\beta p$, one can see that $(\partial\kappa/\partial\beta p)_{\beta}=0$ yields
$\Omega^{\prime}(-\kappa+\beta p,\beta)=\Omega^{\prime}(\beta p,\beta).$ (12)
The Widom line can be analytically continued as a branch lying above the FW line by requiring that $\zeta^{-1}$ is maximal, i.e. $(\partial\zeta/\partial\beta p)_{\beta}=0$.
@@
Up to this point, we now have all the necessary ingredients to compute the Zeno, Seno, FW and Widom lines for the one-dimensional Jagla fluid. But before doing that, and for the sake of completeness, we will take advantage of the simple form of the intermolecular potential $\phi(r)$ of this fluid, as given by Equation (3), to obtain explicitly its second virial coefficient. This will provide us with the means to compute also the Boyle temperature. The explicit analytic result for the second virial coefficient reads
@@
Figure 1: Zeno, Seno, FW and Widom curves in the $T^{*}$ vs $\rho^{*}$ plane for a one-dimensional Jagla fluid with the parameters given in Equation (6). The open circle at $\rho^{*}=0$ represents the Boyle temperature $T_{B}^{*}\simeq 0.4758$. Below the Zeno line, one has $Z<1$, while $Z>1$ above it. Similarly, $\chi_{T}>1$ below the Seno line and $\chi_{T}<1$ above it. Furthermore, below the FW line, the decay of $h(r)$ is monotonic, while it is oscillatory above it. The Widom line is the locus of points where the correlation length is maximal at a given temperature. While the Zeno, Seno and FW lines terminate at the Boyle density $\rho_{B}^{*}=\sigma/\lambda_{1}\simeq 0.77$, the Widom line ends at $\rho_{B}^{*}/2\simeq 0.38$.
In Figure 1 we show the resulting Zeno, Seno, FW and Widom lines in the temperature vs density plane. Note that, while the Zeno and Seno lines do start at the Boyle temperature, the FW line diverges for $\rho\to 0$, despite the wrong impression one might get from the figure. Two more things are also worth pointing out at this stage. On the one hand, the Zeno line is not a straight line and ends at the Boyle density $\rho_{B}=\lambda_{1}^{-1}$; the same density is the zero-temperature end of the Seno and FW lines, while the Widom line terminates at $\rho_{B}/2$ (see the Appendix for a proof). On the other hand, it is clear that for this system the conjecture of Stopper et al. [17] concerning the Seno and FW lines is not sustained. Whether it will hold for the three-dimensional Jagla fluid will be discussed in Section 3.
@@ 3.2 Rational-function approximation
In a previous paper [16], some of us presented a semi-analytical approach based on the RFA [32, 33, 25] to obtain $g(r)$, including its asymptotic behaviour for large $r$. The application of the RFA to the Jagla fluid was made by assuming that a discretised version of the potential given in Equation (3) consisting in a hard core plus of a sequence of $n$ steps of heights $\varepsilon_{j}$ and widths $\sigma_{j}-\sigma_{j-1}$ (with the conventions $\sigma_{0}=\sigma$ and $\sigma_{n}=\lambda_{2}$), leads to essentially the same cavity function as the original Jagla potential. By considering the second virial coefficient and some representative cases, it was found that the choice $n=10$ proved to be a reasonable one, leading to good agreement with MC simulation results. Such an agreement worsened as the density increased and/or the temperature decreased, especially near contact. But, even in those cases, the oscillations of $g(r)$ for larger distances were well accounted for, at least at a qualitative level.
@@ 3.3 Monte Carlo simulations
For the FW line, the main problem is how to know from the MC data of $g(r)$ at a given state $(\rho^{*},T^{*})$ sufficiently close to the line whether that state is above the line (region of oscillatory decay) or below it (region of monotonic decay). If $g(r)$ were known with a good signal-to-noise ratio in the asymptotic large-$r$ domain, it would be in principle possible to assess whether the decay is oscillatory or monotonic since one of the two competing behaviours in Equation (18) would dominate. However, the closer the state is to the line, the closer the values of $\kappa$ and $\zeta$ become. Consequently, larger distances are required to observe the prevalence of one of the two competing behaviours. In addition, it is worth noting that the amplitude $A_{\kappa}$ of the monotonic behaviour is typically smaller than the amplitude $2|A_{\zeta}|$ of the oscillatory behaviour. As a result, the oscillatory behaviour can overshadow the monotonic behaviour for intermediate distances, even if $\kappa<\zeta$, a feature that was also observed and reported by Stopper et al. [37] for patchy particles. We have also noted that the fitting in Equation (20), although suitable for measuring $\chi_{T}$, lacks robustness in determining whether $\kappa<\zeta$ or $\kappa>\zeta$.
@@
Figure 2 shows $r|h(r)|$ (in logarithmic scale), as obtained from the RFA, for a density $\rho^{*}=0.30$ and four temperatures: $T^{*}=0.60$, $0.65$, $0.70$ and $0.75$. For this density, the RFA temperature corresponding to the FW line is known to be $T^{*}=0.7315$. The first thing to note in this case is that the approximation with the two leading poles is able to capture the whole total correlation function for distances beyond $r\simeq 3\sigma$ for all temperatures. Next, we note that the signature that one is sufficiently below the temperature corresponding to the FW line is that ‘anomalous’ neighbouring nodes appear [cf. Figures 2(a–c)]. These nodes, which eventually disappear for large enough distances [although oscillations may still be seen, cf. Figure 2(a)], exhibit an anomalous behavior: their separation is smaller than that of neighbouring nodes and their maxima always fall below that of the neighbouring peaks [cf. Figure 2(a)]. As the temperature increases, remaining below the FW line, the anomalous nodes become progressively less apparent within the range $r<10\sigma$ [cf. Figures 2(b,c)]. Finally, when one is close to or above the temperature corresponding to the FW line, the nodes become regular [cf. Figure 2(d)]. According to our criterion, one would conclude that $T^{*}=0.65$, or even $T^{*}=0.70$, are lower-bound estimates for the temperature of the FW line when $\rho^{*}=0.30$, which agrees with the true FW temperature $T^{*}=0.7315$ predicted by the RFA for $\rho^{*}=0.30$.
@@
Figure 4: Zeno, Seno, FW and Widom curves in the $T^{*}$ vs $\rho^{*}$ plane for a three-dimensional Jagla fluid with the parameters given in Equation (6). The lines are RFA predictions and the symbols represent estimates obtained from our MC simulations. The open circle at $\rho^{*}=0$ represents the Boyle temperature $T_{B}^{*}\simeq 1.3879$. Below the Zeno line, one has $Z<1$, while $Z>1$ above it. Similarly, $\chi_{T}>1$ below the Seno line and $\chi_{T}<1$ above it. Furthermore, below the FW line, the decay of $h(r)$ is monotonic, while it is oscillatory above it. The Widom line is the locus of points where the correlation length is maximal at a given temperature.
@@ 3.4 Results
In order to set the proper perspective for the assessment of our findings, in Figure 4 we show the resulting Zeno, Seno and FW lines for the three-dimensional Jagla fluid, as obtained both from the RFA approach (with a discretisation of $n=10$ steps) and from simulation. The Widom line predicted by the RFA [16] is also included. It terminates at the critical point $(\rho_{c}^{*},T_{c}^{*})=(0.162,0.574)$, which slightly shifts to $(\rho_{c}^{*},T_{c}^{*})=(0.160,0.577)$ if $n=20$ is employed.
@@
Figure 5: Plot of (a) $g(r)$ and (b) $r|h(r)|$ (in logarithmic scale) for a three-dimensional Jagla fluid with the parameters given in Equation (6) and $\rho^{*}=0.30$, $T^{*}=0.75$. The solid lines correspond to the RFA, while the dashed lines represent MC simulation data. Note that the hard-core diameter $\sigma=1$ has been taken as the unit of length.
To illustrate how the discrepancies between the RFA and MC simulations for the transition lines are consistent with a reasonable global agreement in the radial distribution function, we compare the RFA and MC values of $g(r)$ and $r|h(r)|$ at a density $\rho^{*}=0.30$ and a temperature $T^{*}=0.75$ in Figure 5. These conditions correspond to the scenarios depicted in Figures 2(d) and 3(a), respectively. A remarkable overall agreement is observed, although the RFA tends to slightly underestimate $g(r)$ within the interval $\sigma\leq r\leq\lambda_{1}$ and near the second maximum. Considering Equations (2) and (16), this suggests that the RFA tends to underestimate the values of $Z$ and $\chi_{T}$. Consequently, this leads to an upward shift of the Zeno line and a downward shift of the Seno line with respect to the MC values. It is also evident from Figure 5(a) that $g(r)\simeq 1$ for $r>4\sigma$ if $\rho^{*}=0.30$ and $T^{*}=0.75$. This makes it rather challenging to determine whether the asymptotic decay is monotonic or oscillatory. In the case of the RFA, we know from the pole analysis of the Laplace transform $\mathcal{G}(s)$ that the decay is oscillatory, while our criterion suggests that the decay of the MC data is monotonic. This distinction is clearly apparent in Figure 5(b).
The previous observations indicate that, in view of its already known limitations, the good qualitative (and even quantitative) performance of the RFA approach observed in a certain region of the phase diagram may be lost under very stringent conditions of high density and low temperature.